OpenStudy (anonymous):

FELLOW SMART PEOPLE! what is the answer to x2-11x+19=-5

OpenStudy (anonymous):

do the quadratic formula

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

oh! that's right! the quad.. duh.. thanks

OpenStudy (anonymous):

x=3, I solved it mentally

OpenStudy (anonymous):

then maybe I should check it

OpenStudy (anonymous):

oh wait. the paper says solve it by factoring.

OpenStudy (anonymous):

3^2=9 11*3=33 9-33=-24 -24+19=-5 Now can I get a medal?

OpenStudy (anonymous):

gotta factor it yo.